# Scaffold Diffusion: Sparse Multi-Category Voxel Structure Generation with Discrete Diffusion

### v1

@@ Abstract
Machine Learning, ICML
footnotetext: Preprint version. 1 Chan Zuckerberg Imaging Institute.
[…]
@@ 2 Related Work
Recent advances have shown the promise of 3D generative models. PointVoxelDiffusion (Zhou et al., 2021) demonstrates the effectiveness of diffusion models for point cloud generation using point-voxel representations. LiON (Zeng et al., 2022) introduces a hierarchical latent diffusion model capable of both point cloud and mesh generation. For large scale generation, XCube (Ren et al., 2024) presents a hierarchical latent diffusion model for sparse binary-valued voxel maps; for textured structures they apply an off the shelf texture model post-generation. In the domain of floor plan generation, MiDiffusion (Hu et al., 2024) introduces a mixed continuous-discrete diffusion model that simultaneously predicts categorical labels and geometric features.
Most relevant to our work, Lee et al. (Lee et al., 2023) pioneer the application of discrete diffusion models for 3D categorical data. They apply multinomial diffusion (Hoogeboom et al., 2021) to 3D segmentation map generation, and develop both a standalone multinomial diffusion model which simply operates on an entire segmentation map and a two stage VQ-VAE latent multinomial diffusion model which operates on a smaller latent.
@@ 2.2 Minecraft Generative Models
VoxelCNN (Chen et al., 2019) introduces the 3D-Craft Dataset of human created minecraft houses and develops an autoregressive 3D convolution network that predicts next block placements given previous block placements. Their model requires previously placed blocks as input and has not been evaluated on complete structure generation, only partial generation.
WorldGAN (Awiszus et al., 2021) attempts large-scale Minecraft world generation using a GAN-based approach with word2vec-like tokenization scheme. However, as the authors acknowledge, their method fails to generate functionally coherent structures such as houses. Alternately, DreamCraft (Earle et al., 2024) employs a NeRF-based model and Sudhakaran et al. (Sudhakaran et al., 2021) uses a neural cellular automata based approach. Both methods however require re-training for each sample, with DreamCraft requiring hours to generate a single sample.
@@ 2.3 Discrete Diffusion Models
Sohl-Dickstein et al. (Sohl-Dickstein et al., 2015) first introduces diffusion for discrete spaces with a diffusion process over binary random variables. This work was extended to categorical variables by Multinomial Diffusion (Hoogeboom et al., 2021) with a uniform transition process. D3PM (Austin et al., 2021) generalizes Multinomial Diffusion and establishes a discrete diffusion framework with arbitrary transition matrices.
While discrete diffusion has shown promise for generating discrete data, they have generally been considered to have inferior generation quality compared to their autoregressive counterparts. Recent work such as MDLM (Sahoo et al., 2024) has narrowed the performance gap between discrete diffusion and autoregressive models for text generation through a simplified continuous time ELBO objective, achieving improved sample quality and computational efficiency.
@@ 3 Preliminary
Diffusion models have been shown to be effective generative models for many data distributions. While diffusion models were initially popularized for continuous data generation, such as image synthesis (Ho et al., 2020), discrete diffusion models have been extended to discrete data, such as text (Austin et al., 2021). D3PM (Austin et al., 2021) introduces a general framework of discrete diffusion models and evaluates on different forward transition matrices, such as uniform, absorbing state, and discretized gaussian transition matrices. Masked Discrete Language Model (MDLM) is an absorbing state variant of discrete diffusion that has shown effective performance under a simplified training objective (Sahoo et al., 2024). Formally, MDLM operates on discrete token sequences $\mathbf{x}=(x_{1},x_{2},\ldots,x_{L})$ where each token $x_{i}$ belongs to a finite vocabulary $\mathcal{V}$. The sequence $\mathbf{x}\in\mathbb{Z}^{L}$ evolves to a sequence of hidden latents $\mathbf{z_{t}}$ according to the forward corrupting Markov chain defined by the absorbing state transition kernel matrix $Q_{t}$, where the discrete time forward transition is $q(z_{t}|z_{t-1})=Cat(z_{t};z_{t-1}Q_{t})$. The absorbing state transition kernel matrix $Q_{t}$ is defined such that each token transitions to the absorbing $[MASK]$ token with probability $\beta_{t}$ and remains the same with probability $1-\beta_{t}$. The marginal of the forward process is defined as $q(\mathbf{x_{t}}|\mathbf{x})=Cat(\mathbf{x_{t}};\alpha_{t}\mathbf{x}+(1-\alpha_{t})\mathbf{m})$, where $\mathbf{m}$ is the Dirac-delta distribution on the mask token.
@@ 3.1 Training Objective
Figure 2: Sample generation quality comparison between Scaffold Diffusion and baselines; Scaffold Diffusion (top row), autoregressive baseline (middle row), and (Lee et al., 2023) (bottom row). While Scaffold Diffusion can generate realistic and functional 3D structures, the autoregressive baseline generates structures dominated by a few block types or structures with implausible block placements. (Lee et al., 2023) suffers from an over-representation of background voxels.
@@ 4 Method
Because of the memory and class imbalance challenges of working with the full voxel map $\mathbf{O}\in\mathbb{Z}^{D\times D\times D}$, we extract only the $k$ many occupied voxels and their corresponding voxel location $\{(x_{i},y_{i},z_{i})\}_{i=1}^{k}$. From these occupied voxels, we extract their locations to define a $L\geq k$ length sequence $\mathbf{x}\in\mathbb{Z}^{L}$ (potentially appended with a variable number of padding tokens) for the masked diffusion language model to generate. For our model to be spatially aware during its generation process, we integrate 3D positional embedding. We find 3D sinusoidal positional embeddings as in (Mo et al., 2023) effective. Finally, given a generated sequence $\mathbf{x}_{0}\in\mathbb{Z}^{L}$ and the occupied voxel locations, we can re-construct a generated voxel map $\mathbf{X}\in\mathbb{Z}^{D\times D\times D}$ for visualization purposes.
@@ 5 Experiments
We evaluate on the 3D-Craft dataset (Chen et al., 2019), a dataset of human-created Minecraft houses. The 3D-Craft dataset contains a time-stamped sequence of block placements $\{(x_{t},y_{t},z_{t},id_{t})\}_{t}$. The Minecraft block IDs are integers in the range $[0,255]$ and we have in our dataset $n=253$ possible block ID values. Since in many settings a sequence of block placements is not available, we convert the sequences to a voxel structure by placing the block placements in a voxel cube of some pre-defined dimension $\mathbf{X}\in\mathbb{Z}^{D\times D\times D}$.
@@ 5.1 Implementation Details
For our discrete diffusion model, we use the Masked Diffusion Language Model MDLM (Sahoo et al., 2024). We adopt their code and use their model architecture, optimizer, and hyperparameters. Specifically, we use the Diffusion Transformer (DiT) backbone (Peebles & Xie, 2023) with $n=12$ blocks and $n=12$ heads and a sequence length of $L=1024$. We use a log-linear noise schedule and for faster inference we use cached updates. We use EMA with $\beta=0.9999$ and AdamW optimizer with a learning rate of $lr=3\times 10^{-4}$, $\beta_{1}=0.9$, $\beta_{2}=0.999$ and $\epsilon=1\times 10^{-8}$ and a constant warm-up of $2500$ steps. We train our models with a maximum of $10^{6}$ steps.
We additionally integrate 3D sinusoidal positional embeddings as in (Mo et al., 2023) into our DiT backbone.
@@ 5.2 Baselines
As a comparable baseline, we implement a VQ-VAE latent multinomial diffusion model as in (Lee et al., 2023). In line with their method, the model is trained with an inverse class-frequency weighted cross-entropy loss to mitigate the effects of class imbalance. We adopt their default hyperparameter settings and their provided code.
@@ 5.3 Qualitative Results
While previous works such as VoxelCNN report quantitative metrics such as perplexity and next-action accuracy in addition to qualitative results, we recognize that evaluation of generative capability is inherently a qualitative task. In Figure 2 we demonstrate the generative capability of Scaffold Diffusion in comparison to (Lee et al., 2023) and an autoregressive version of the model. We note that Scaffold Diffusion can generate spatially consistent and functional structures, whereas the autoregressive version typically generates collapsed structures that contain only a few different block categories types. (Lee et al., 2023) suffers from an overrepresentation of background tokens, which we hypothesize is due to its training on the entire voxel structure, even with inverse class-frequency loss re-weighting and VQ-VAE latent diffusion.
@@ 5.4 Ablations
We ablate our backbone and design decision to operate on a sequence of occupied token positions and consider operating on the entire voxel map, similar to (Lee et al., 2023). We adopt DiT-3D (Mo et al., 2023) which voxelizes point clouds into voxel patches and use our voxel map directly to construct patches for the DiT transformer. We choose their default patch size of $p=4$ and use their DiT model with a depth of $n=12$ and $n=12$ heads. Similar to the behavior with (Lee et al., 2023), even with inverse class-frequency loss re-weighting, we suffer from the problem of sparsity and background voxel class imbalance and fail to generate plausible boolean occupancy structures.

### v3

@@ 2 Related Work
Recent advances have shown the promise of 3D generative models. PointVoxelDiffusion [zhou2021pvd] demonstrates the effectiveness of diffusion models for point cloud generation using point-voxel representations. LiON [zeng2022lion] introduces a hierarchical latent diffusion model capable of both point cloud and mesh generation. For large scale generation, XCube [ren2024xcube] presents a hierarchical latent diffusion model for sparse binary-valued voxel maps; for textured structures they apply an off the shelf texture model post-generation. In the domain of floor plan generation, MiDiffusion [hu2024midiffusion] introduces a mixed continuous-discrete diffusion model that simultaneously predicts categorical labels and geometric features.
Most relevant to our work, Lee et al. [lee2023diffusion] pioneer the application of discrete diffusion models for 3D categorical data. They apply multinomial diffusion [hoogeboom2021argmax] to 3D segmentation map generation, and develop both a standalone multinomial diffusion model which simply operates on an entire segmentation map and a two stage VQ-VAE latent multinomial diffusion model which operates on a smaller latent.
@@ 2.2 Minecraft Generative Models
VoxelCNN [chen2019voxelcnn] introduces the 3D-Craft dataset of human created minecraft houses and develops an autoregressive 3D convolution network that predicts next block placements given previous block placements. Their model requires previously placed blocks as input and has not been evaluated on complete structure generation, only partial generation.
WorldGAN [awiszus2021worldgan] attempts large-scale Minecraft world generation using a GAN-based approach with word2vec-like tokenization scheme. However, as the authors acknowledge, their method fails to generate functionally coherent structures such as houses. Alternately, DreamCraft [earle2024dreamcraft] employs a NeRF-based model and Sudhakaran et al. [sudhakaran2021growing] uses a neural cellular automata based approach. Both methods however require re-training for each sample, with DreamCraft requiring hours to generate a single sample.
Oasis [oasis2024] develops a next-frame autoregressive Minecraft world model that generates action-conditioned future frames, allowing the user to interact with the simulated game environment.
@@ 2.3 Discrete Diffusion Models
Sohl-Dickstein et al. [sohldickstein2015deep] first introduces diffusion for discrete spaces with a diffusion process over binary random variables. This work was extended to categorical variables by Multinomial Diffusion [hoogeboom2021argmax] with a uniform transition process. D3PM [austin2021d3pm] generalizes Multinomial Diffusion and establishes a discrete diffusion framework with arbitrary transition matrices.
While discrete diffusion has shown promise for generating discrete data, they have generally been considered to have inferior generation quality compared to their autoregressive counterparts. Recent work such as MDLM [sahoo2024mdlm, shi2025simplifiedgeneralizedmaskeddiffusion] has narrowed the performance gap between discrete diffusion and autoregressive models for text generation through a simplified continuous time ELBO objective, achieving improved sample quality and computational efficiency.
@@ 3 Preliminary
Diffusion models have been shown to be effective generative models for many data distributions. While diffusion models were initially popularized for continuous data generation, such as image synthesis [ho2020denoising], discrete diffusion models have been extended to discrete data, such as text [austin2021d3pm]. D3PM [austin2021d3pm] introduces a general framework of discrete diffusion models and evaluates on different forward transition matrices, such as uniform, absorbing state, and discretized gaussian transition matrices. Masked Discrete Language Model (MDLM) is an absorbing state variant of discrete diffusion that has shown effective performance under a simplified training objective [sahoo2024mdlm]. Formally, MDLM operates on discrete token sequences $\mathbf{x}=(x_{1},x_{2},\ldots,x_{L})$ where each token $x_{i}$ belongs to a finite vocabulary $\mathcal{V}$. The sequence $\mathbf{x}\in\mathbb{Z}^{L}$ evolves to a sequence of hidden latents $\mathbf{z_{t}}$ according to the forward corrupting Markov chain defined by the absorbing state transition kernel matrix $Q_{t}$, where the discrete time forward transition is $q(z_{t}|z_{t-1})=Cat(z_{t};z_{t-1}Q_{t})$. The absorbing state transition kernel matrix $Q_{t}$ is defined such that each token transitions to the absorbing $[MASK]$ token with probability $\beta_{t}$ and remains the same with probability $1-\beta_{t}$. The marginal of the forward process is defined as $q(\mathbf{x_{t}}|\mathbf{x})=Cat(\mathbf{x_{t}};\alpha_{t}\mathbf{x}+(1-\alpha_{t})\mathbf{m})$, where $\mathbf{m}$ is the Dirac-delta distribution on the mask token.
@@ 3.1 Training Objective
Figure 2: Sample generation quality comparison between Scaffold Diffusion and baselines; Scaffold Diffusion (top row), autoregressive baseline (middle row), and [lee2023diffusion] (bottom row). While Scaffold Diffusion can generate realistic and functional 3D structures, the autoregressive baseline generates structures dominated by a few block types or structures with implausible block placements. [lee2023diffusion] suffers from an over-representation of background voxels.
@@ 4 Method
Because of the memory and class imbalance challenges of working with the full voxel map $\mathbf{O}\in\mathbb{Z}^{D\times D\times D}$, we extract only the $k$ many occupied voxels and their corresponding voxel location $\{(x_{i},y_{i},z_{i})\}_{i=1}^{k}$. From these occupied voxels, we extract their locations to define a $L\geq k$ length sequence $\mathbf{x}\in\mathbb{Z}^{L}$ (potentially appended with a variable number of padding tokens) for the masked diffusion language model to generate. For our model to be spatially aware during its generation process, we integrate 3D positional embedding. We find 3D sinusoidal positional embeddings as in [mo2023dit] effective. Finally, given a generated sequence $\mathbf{x}_{0}\in\mathbb{Z}^{L}$ and the occupied voxel locations, we can re-construct a generated voxel map $\mathbf{X}\in\mathbb{Z}^{D\times D\times D}$ for visualization purposes.
@@ 5 Experiments
We evaluate on the 3D-Craft dataset [chen2019voxelcnn], a dataset of human-created Minecraft houses. The 3D-Craft dataset contains a time-stamped sequence of block placements $\{(x_{t},y_{t},z_{t},id_{t})\}_{t}$. The Minecraft block IDs are integers in the range $[0,255]$ and we have in our dataset $n=253$ possible block ID values, leading to a vocabulary size of $|V|=253$. Since in many settings a sequence of block placements is not available, we convert the sequences to a voxel structure by placing the block placements in a voxel cube of some pre-defined dimension $\mathbf{X}\in\mathbb{Z}^{D\times D\times D}$.
@@ 5.1 Implementation Details
For our discrete diffusion model, we use the Masked Diffusion Language Model MDLM [sahoo2024mdlm]. We adopt their code and use their model architecture, optimizer, and hyperparameters. Specifically, we use the Diffusion Transformer (DiT) backbone [peebles2023scalable] with $n=12$ blocks and $n=12$ heads and a sequence length of $L=1024$. We use a log-linear noise schedule and for faster inference we use cached updates. We use EMA with $\beta=0.9999$ and AdamW optimizer with a learning rate of $lr=3\times 10^{-4}$, $\beta_{1}=0.9$, $\beta_{2}=0.999$ and $\epsilon=1\times 10^{-8}$ and a constant warm-up of $2500$ steps. We train our models with a maximum of $10^{6}$ steps. We additionally integrate 3D sinusoidal positional embeddings as in [mo2023dit] into our DiT backbone.
All experiments were trained using a single RTX 5090 and each experiment took under 12 hours to finish training.
@@ 5.2 Baselines
As a comparable baseline, we implement a VQ-VAE latent multinomial diffusion model as in [lee2023diffusion]. In line with their method, the model is trained with an inverse class-frequency weighted cross-entropy loss to mitigate the effects of class imbalance. We adopt their default hyperparameter settings and their provided code.
@@
Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption
Figure 3: Diversity of generated samples. Scaffold Diffusion produces varied and realistic 3D structures for the same occupancy map.
@@ 5.3 Qualitative Results
While previous works such as VoxelCNN report quantitative metrics such as perplexity and next-action accuracy in addition to qualitative results, we recognize that evaluation of generative capability is inherently a qualitative task. In Figure 2 we demonstrate the generative capability of Scaffold Diffusion in comparison to [lee2023diffusion] and an autoregressive version of the model. We note that Scaffold Diffusion can generate spatially consistent and functional structures, whereas the autoregressive version typically generates collapsed structures that contain only a few different block categories types. [lee2023diffusion] suffers from an overrepresentation of background tokens, which we hypothesize is due to its training on the entire voxel structure, even with inverse class-frequency loss re-weighting and VQ-VAE latent diffusion.
Scaffold Diffusion also demonstrates diverse generative capability. In Figure 3 we illustrate the diversity of structures that Scaffold Diffusion generates for the same conditional occupancy map.
@@ 5.4 Ablations
We ablate our backbone and design decision to operate on a sequence of occupied token positions and consider operating on the entire voxel map, similar to [lee2023diffusion]. We adopt DiT-3D [mo2023dit] which voxelizes point clouds into voxel patches and use our voxel map directly to construct patches for the DiT transformer. We choose their default patch size of $p=4$ and use their DiT model with a depth of $n=12$ and $n=12$ heads. Similar to the behavior with [lee2023diffusion], even with inverse class-frequency loss re-weighting, we suffer from the problem of sparsity and background voxel class imbalance and fail to generate plausible boolean occupancy structures.